A contractor wants to use 34 feet of molding, cut into three pieces, to trim the sides and the top of a garage door. The long piece is 1.5 feet longer than three times the length of each shorter piece. Find the length of each piece.
Please show a step by step process.

Answer by rfer(16322)

3x+1.5+2x=34
5x=34-1.5
5x=32.5
x=32.5/5
x=6.5ft two sides each
3x+1.5=21ft top